Waterways Problem 05

The Vice President for Sales and Marketing at Waterways Corporation is planning for production needs to meet sales demand in the coming year. He is also trying to determine how the company’s profits might be increased in the coming year. This problem asks you to use cost-volume-profit concepts to help Waterways understand contribution margins of some of its products and decide whether to mass-produce any of them.

Waterways markets a simple water control and timer that it mass-produces. Last year, the company sold 751,000 units at an average selling price of $4.90 per unit. The variable costs were $2,575,930, and the fixed costs were $772,779.

What is the product’s contribution margin ratio? (Round ratio to 0 decimal places, e.g. 25%.)
Contribution margin ratio %
Answer #1

Total sales=751000*4.9=$3679900

Contribution margin=Sales-Variable cost  

=3679900-2,575,930=$1103970

Contribution margin ratio=Contribution margin/Sales

=1103970/3679900

=30%